# Edit Distance Robust Watermarks for Language Models

### v1

@@ 3 Secret-key substitution PRCs from weaker assumptions
In this section, we discuss a new construction of binary PRCs for substitution channels. Though such PRCs were also obtained by [CG24], the codes in [CG24] relied on relatively strong average-case hardness assumptions, in the sense that they imply the existence of public-key cryptography (i.e., in the context of Impagliazzo’s Five Worlds [Imp95], they imply primitives in “Cryptomania”). In contrast, our construction relies only on the hardness of the existence of a family of pseudorandom functions that enjoys a certain locality property; such an assumption is generally believed to be weaker than the ones in [CG24], in the sense that it is only known to yield cryptographic primitives in “Minicrypt”.
@@
Fix functions $\ell(\lambda),n(\lambda):\mathbb{N}\to\mathbb{N}$ of a security parameter $\lambda$, and a collection of functions $\{F_{s}:\{0,1\}^{n(\lambda)}\to\{0,1\}\}$, indexed by $s\in\{0,1\}^{\ell(\lambda)}$, for $\lambda\in\mathbb{N}$. We say that the collection $\{F_{s}\}_{s}$ is a weak pseudorandom function family (weak PRF) if for every probabilistic polynomial-time algorithm $\mathsf{Adv}$, it holds that
$\left|\mathbb{E}_{s\sim\{0,1\}^{n(\lambda)}}\left[\widetilde{\mathsf{Adv}}^{F_{s}(\cdot)}(1^{n(\lambda)})=1\right]-\mathbb{E}_{F_{\mathsf{Unif}}}\left[\widetilde{\mathsf{Adv}}^{F_{\mathsf{Unif}}}(1^{n(\lambda)})\right]\right|\leq\mathsf{negl}(\lambda),$ (2)
where $\widetilde{\mathsf{Adv}}^{G}$, for a mapping $G:\{0,1\}^{n(\lambda)}\to\{0,1\}$ means that $\mathsf{Adv}$ can make calls to the function $G$, and for each call receives a tuple $(x,G)$, for $x\sim\operatorname{Unif}(\{0,1\}^{n})$. (In particular, the tilde refers to the fact that $\mathsf{Adv}$ can only call $G(x)$ on a uniformly chosen $x$.) Moreover, ${F_{\mathsf{Unif}}}:\{0,1\}^{n}\to\{0,1\}$ denotes a uniformly random function. We will often refer to the function $G$ that $\mathsf{Adv}$ can make queries to as the oracle $\mathsf{Adv}$ has access to. Given $q\in[0,1/2)$, we say that $\{F_{s}\}_{s}$ is a weak PRF family with noise level $q$ if Eq. 2 holds where each call to $F_{s}(\cdot)$ by $\mathsf{Adv}$ returns $(x,F_{s}(x)\oplus e)$ where $x\sim\operatorname{Unif}(\{0,1\})^{n}$ and $e\sim\mathrm{Ber}(q)$.
@@
  • $\mathsf{KeyGen}(1^{\lambda})$ chooses a function in $\mathcal{F}$ (indexed by $s$) together with a uniformly random $z\in\operatorname{Unif}(\{0,1\}^{N(\lambda)})$ and a uniform permutation $\pi:[N(\lambda)]\to[N(\lambda)]$, and returns $\mathsf{sk}=(s,z,\pi)$.
  • $\mathsf{Encode}(1^{\lambda},(s,z,\pi),\emptyset)$ draws $m(\lambda)$ uniformly random elements of $\{0,1\}^{n(\lambda)}$, $x_{1},\ldots,x_{m(\lambda)}$, applies $F_{s}$ to each of them and flips the result with probability $q$ to obtain bits $(w_{1},\ldots,w_{m(\lambda)})$, and perturbs the concatenation $((x_{i},w_{i}))_{i\in[m(\lambda)]}$ according to $z,\pi$ as on 9.
@@
8:     Define $a:=((x_{1},F_{s}(x_{1})\oplus e_{1}),\ldots,(x_{m(\lambda)},F_{s}(x_{m(\lambda)})\oplus e_{m(\lambda)}))$.
9:     return the string $\pi\circ(a\oplus z)$, where $a$ is padded with $N(\lambda)-(n(\lambda)+1)m(\lambda)$ zeros.
10:function $\mathsf{Decode}$($1^{\lambda},\mathsf{sk}=(s,z,\pi),y$)
11:     Let $x_{1},\ldots,x_{m(\lambda)}\in\{0,1\}^{n(\lambda)}$ and $w_{1},\ldots,w_{m(\lambda)}\in\{0,1\}$ be defined by $((x_{1},w_{1}),\ldots,(x_{m(\lambda)},w_{m(\lambda)}))=((\pi^{-1}\circ y)\oplus z)_{1:(n(\lambda)+1)m(\lambda)}$.
12:     Let $W:=\sum_{j=1}^{m(\lambda)}\mathbbm{1}\{{w_{j}=F_{s}(x_{j})}\}$.
13:     If $W>\frac{m(\lambda)}{2}+\ln(m(\lambda))\sqrt{m(\lambda)}$, return $\emptyset$; otherwise, return $\perp$.
@@
The bulk of the proof lies in establishing robustness to $p$-bounded substitution channels (Lemma A.8). Let us first make sure that $\mathsf{Decode}(1^{\lambda},\mathsf{sk},y)$ returns $\emptyset$ with high probability if its input $y$ is simply an output of $\mathsf{Encode}(1^{\lambda},\mathsf{sk},\emptyset)$ (i.e., if the channel $\mathcal{E}$ does nothing to its input). Indeed, the tuples $(x_{i},w_{i})$ (for $i\in[m(\lambda)]$) computed in 11 of $\mathsf{Decode}$ are exactly equal to $(x_{i},F_{s}(x_{i})\oplus e_{i})$. As long as the noise rate $q$ of $e_{i}$ is bounded away from $1/2$, the statistic $W$ computed in $\mathsf{Decode}$ will be larger than $m(\lambda)/2+\ln(m(\lambda))\sqrt{m(\lambda)}$ with all but negligible probability, as desired.
Now what if $y$ is drawn from $\mathcal{E}(y^{0})$, for $y^{0}\leftarrow\mathsf{Encode}(1^{\lambda},\mathsf{sk},\emptyset)$, for some $p$-bounded substitution channel $\mathcal{E}$? Then the tuples computed in 11 of $\mathsf{Decode}$ may be written as $(x_{i}^{\prime},w_{i}^{\prime})$ (for $i\in[m(\lambda)]$), where $(x_{i}^{\prime},w_{i}^{\prime})$ are “perturbed” versions of $(x_{i},w_{i})$, where $x_{i}\sim\operatorname{Unif}(\{0,1\}^{n(\lambda)}),w_{i}=F_{s}(x_{i})\oplus e_{i}$ are as computed in $\mathsf{Encode}$. Although the channel $\mathcal{E}$ may introduce substitutions in an adversarial fashion (i.e., it may not introduce substitutions at each position independently with probability $p$), by virtue of the fact that the output string of $\mathsf{Encode}$ is $\pi\circ(a\oplus z)$ for uniformly random $z\in\{0,1\}^{N(\lambda)},\pi:[N(\lambda)]\to[N(\lambda)]$, we can show that $(x_{i}^{\prime},w_{i}^{\prime})$ is close to being distributed by perturbing each bit of $(x_{i},w_{i})$ independently with some probability $p^{\prime}\leq p$. The proof of this fact uses an approximation of a Hypergeometric distribution with a Binomial distribution: roughly speaking, the permutation $\pi$ allows us to “pick out”, for each $i$, a uniformly random subset of $n(\lambda)+1$ positions corresponding to $(x_{i}^{\prime},w_{i}^{\prime})$, out of $N(\lambda)$ total positions of the string $y$. Of these, $p^{\prime}\cdot N(\lambda)\leq p\cdot N(\lambda)$ are changed by $\mathcal{E}$, and as long as $N(\lambda)\gg n(\lambda)$, applying $\mathcal{E}$ thus has nearly the same effect as changing each of the $n(\lambda)+1$ positions of $(x_{i},w_{i})$ independently with probability $p^{\prime}$.
@@ A.2 Additional preliminaries
for some constant $c>0$, then we could construct an adversary $\mathsf{Adv}^{\prime}$ which violates security of $F_{s}$ (namely, Definition 3.1), as follows: $\mathsf{Adv}^{\prime}$ makes $Qm$ calls to $F_{s}(\cdot)$ on random inputs, which gives it samples $(x_{j}^{(i)},F_{s}(x_{j}^{(i)})\oplus e_{j}^{(i)})$ for $j\in[m],i\in[Q]$, draws a uniformly random permutation $\pi:[N]\to[N]$ and a uniformly random string $z\in\{0,1\}^{N}$, and outputs the result of $\mathsf{Adv}$ given the $Q$ tuples $y^{(i)}:=\pi\circ(((x_{1}^{(i)},F_{s}(x_{1}^{(i)})\oplus e_{1}^{(i)}),\ldots,(x_{m}^{(i)},F_{s}(x_{m}^{(i)})\oplus e_{m}^{(i)}))\oplus z)$ (for $i\in[Q]$). Note that if the $mQ$ calls made by $\mathsf{Adv}^{\prime}$ were generated by a random function, then the strings $y^{(i)}$ are uniformly random in $\{0,1\}^{N}$, conditioned on the event that $x_{j}^{(i)}$ are all distinct, for $j\in[m],i\in[Q]$. The probability that all $x_{j}^{(i)}$ are distinct is at least $1-mQ/2^{n}\geq 1-\mathsf{negl}(\lambda)$, meaning that, by Eq. 5, $\mathsf{Adv}^{\prime}$ achieves advantage $\lambda^{-c}-\mathsf{negl}(\lambda)$ on distinguishing a random element of the family $\{F_{s}\}_{s}$ from a uniformly random function, thus contradicting the security of $\mathcal{F}$. ∎
@@
Fix $\lambda\in\mathbb{N}$, and write $n=n(\lambda),\ell=\ell(\lambda),m=m(\lambda),N=N(\lambda)$. Let $\mathcal{E}$ be a $p$-weakly-substitution-bounded channel over $\{0,1\}$. Given $x_{j},e_{j}$ ($j\in[m]$) drawn as in 7 of Algorithm 1, consider the string $a:=((x_{1},F_{s}(x_{1})\oplus e_{1}),\ldots,(x_{m},F_{s}(x_{m})\oplus e_{m}))$, and note that the output of $\mathsf{Encode}(1^{\lambda},\mathsf{sk},\emptyset)$ given that $x_{1},\ldots,x_{m},e_{1},\ldots,e_{m}$ are sampled is $\bar{y}:=\pi\circ(a\oplus z)$. (Note that we abuse notation slightly by padding $a$ with $N-(n+1)m$ zeros so that $a\oplus z\in\{0,1\}^{N}$ is well-defined.) Given $v\in\{0,1\}^{n}$, we let, with slight abuse of notation, $\mathcal{E}(v)\in\{0,1\}^{n}$ denote a random variable drawn from the eponymous distribution $\mathcal{E}(v)\in\Delta(\{0,1\}^{n})$. We have

### v2

# Edit Distance Robust Watermarks via Indexing Pseudorandom Codes

@@ 3 Secret-key substitution PRCs from weaker assumptions
In this section, we discuss a new construction of binary PRCs for substitution channels. Though such PRCs were also obtained by [CG24], the codes in [CG24] relied on relatively strong average-case hardness assumptions, in the sense that they imply the existence of public-key cryptography (i.e., in the context of Impagliazzo’s Five Worlds [Imp95], they imply primitives in “Cryptomania”). In contrast, our construction relies only on the hardness of the existence of a family of pseudorandom functions that enjoys a certain locality property; such an assumption is generally believed to be weaker than the ones in [CG24], in the sense that it is only known to yield cryptographic primitives in “Minicrypt”, though we are not aware of a formal separation.
@@
Fix functions $\ell(\lambda),n(\lambda):\mathbb{N}\to\mathbb{N}$ of a security parameter $\lambda$, and a collection of functions $\{F_{s}:\{0,1\}^{n(\lambda)}\to\{0,1\}\}$, indexed by $s\in\{0,1\}^{\ell(\lambda)}$, for $\lambda\in\mathbb{N}$. We say that the collection $\{F_{s}\}_{s}$ is a weak pseudorandom function family (weak PRF) if for every probabilistic polynomial-time algorithm $\mathsf{Adv}$ which outputs a single bit (i.e., 0 or 1), it holds that
$\left|\mathbb{E}_{s\sim\{0,1\}^{\ell(\lambda)}}\left[\widetilde{\mathsf{Adv}}^{F_{s}(\cdot)}(1^{n(\lambda)})\right]-\mathbb{E}_{F_{\mathsf{Unif}}}\left[\widetilde{\mathsf{Adv}}^{F_{\mathsf{Unif}}}(1^{n(\lambda)})\right]\right|\leq\mathsf{negl}(\lambda),$ (2)
where $\widetilde{\mathsf{Adv}}^{G}$, for a mapping $G:\{0,1\}^{n(\lambda)}\to\{0,1\}$ means that $\mathsf{Adv}$ can make calls to the function $G$, and for each call receives a tuple $(x,G)$, for $x\sim\operatorname{Unif}(\{0,1\}^{n})$. (In particular, the tilde refers to the fact that $\mathsf{Adv}$ can only call $G(x)$ on a uniformly chosen $x$.) Moreover, ${F_{\mathsf{Unif}}}:\{0,1\}^{n}\to\{0,1\}$ denotes a uniformly random function. We will often refer to the function $G$ that $\mathsf{Adv}$ can make queries to as the oracle $\mathsf{Adv}$ has access to. Given $q\in[0,1/2)$, we say that $\{F_{s}\}_{s}$ is a weak PRF family with noise level $q$ if Eq. 2 holds where each call to $F_{s}(\cdot)$ by $\mathsf{Adv}$ returns $F_{s}(x)\oplus e$ where $x\sim\operatorname{Unif}(\{0,1\})^{n}$ and $e\sim\mathrm{Ber}(q)$.
@@
  • $\mathsf{KeyGen}(1^{\lambda})$ chooses a function in $\mathcal{F}$ (indexed by $s$) together with a uniformly random $z\sim\operatorname{Unif}(\{0,1\}^{N(\lambda)})$ and a uniform permutation $\pi:[N(\lambda)]\to[N(\lambda)]$, and returns $\mathsf{sk}=(s,z,\pi)$.
  • $\mathsf{Encode}(1^{\lambda},(s,z,\pi),\emptyset)$ draws $m(\lambda)$ uniformly random elements of $\{0,1\}^{n(\lambda)}$, $x_{1},\ldots,x_{m(\lambda)}$, applies $F_{s}$ to each of them and flips the result with probability $q$ to obtain bits $(w_{1},\ldots,w_{m(\lambda)})$, and perturbs the concatenation $((x_{i},w_{i}))_{i\in[m(\lambda)]}$ according to $z,\pi$ as on 10.
@@
8:     Sample $r\sim\operatorname{Unif}(\{0,1\}^{N(\lambda)-(n(\lambda)+1)m(\lambda)})$.
9:     Define $a:=((x_{1},F_{s}(x_{1})\oplus e_{1}),\ldots,(x_{m(\lambda)},F_{s}(x_{m(\lambda)})\oplus e_{m(\lambda)}),r)$.
10:     return the string $\pi\circ(a\oplus z)$.
11:function $\mathsf{Decode}$($1^{\lambda},\mathsf{sk}=(s,z,\pi),y$)
12:     Let $x_{1},\ldots,x_{m(\lambda)}\in\{0,1\}^{n(\lambda)}$ and $w_{1},\ldots,w_{m(\lambda)}\in\{0,1\}$ be defined by $((x_{1},w_{1}),\ldots,(x_{m(\lambda)},w_{m(\lambda)}))=((\pi^{-1}\circ y)\oplus z)_{1:(n(\lambda)+1)m(\lambda)}$.
13:     Let $W:=\sum_{j=1}^{m(\lambda)}\mathbbm{1}\{{w_{j}=F_{s}(x_{j})}\}$.
14:     If $W>\frac{m(\lambda)}{2}+\ln(m(\lambda))\sqrt{m(\lambda)}$, return $\emptyset$; otherwise, return $\perp$.
@@
The bulk of the proof lies in establishing robustness to $p$-bounded substitution channels (Lemma A.8). Let us first make sure that $\mathsf{Decode}(1^{\lambda},\mathsf{sk},y)$ returns $\emptyset$ with high probability if its input $y$ is simply an output of $\mathsf{Encode}(1^{\lambda},\mathsf{sk},\emptyset)$ (i.e., if the channel $\mathcal{E}$ does nothing to its input). Indeed, the tuples $(x_{i},w_{i})$ (for $i\in[m(\lambda)]$) computed in 12 of $\mathsf{Decode}$ are exactly equal to $(x_{i},F_{s}(x_{i})\oplus e_{i})$. As long as the noise rate $q$ of $e_{i}$ is bounded away from $1/2$, the statistic $W$ computed in $\mathsf{Decode}$ will be larger than $m(\lambda)/2+\ln(m(\lambda))\sqrt{m(\lambda)}$ with all but negligible probability, as desired.
Now what if $y$ is drawn from $\mathcal{E}(y^{0})$, for $y^{0}\leftarrow\mathsf{Encode}(1^{\lambda},\mathsf{sk},\emptyset)$, for some $p$-bounded substitution channel $\mathcal{E}$? Then the tuples computed in 12 of $\mathsf{Decode}$ may be written as $(x_{i}^{\prime},w_{i}^{\prime})$ (for $i\in[m(\lambda)]$), where $(x_{i}^{\prime},w_{i}^{\prime})$ are “perturbed” versions of $(x_{i},w_{i})$, where $x_{i}\sim\operatorname{Unif}(\{0,1\}^{n(\lambda)}),w_{i}=F_{s}(x_{i})\oplus e_{i}$ are as computed in $\mathsf{Encode}$. Although the channel $\mathcal{E}$ may introduce substitutions in an adversarial fashion (i.e., it may not introduce substitutions at each position independently with probability $p$), by virtue of the fact that the output string of $\mathsf{Encode}$ is $\pi\circ(a\oplus z)$ for uniformly random $z\in\{0,1\}^{N(\lambda)},\pi:[N(\lambda)]\to[N(\lambda)]$, we can show that $(x_{i}^{\prime},w_{i}^{\prime})$ is close to being distributed by perturbing each bit of $(x_{i},w_{i})$ independently with some probability $p^{\prime}\leq p$. The proof of this fact uses an approximation of a Hypergeometric distribution with a Binomial distribution: roughly speaking, the permutation $\pi$ allows us to “pick out”, for each $i$, a uniformly random subset of $n(\lambda)+1$ positions corresponding to $(x_{i}^{\prime},w_{i}^{\prime})$, out of $N(\lambda)$ total positions of the string $y$. Of these, $p^{\prime}\cdot N(\lambda)\leq p\cdot N(\lambda)$ are changed by $\mathcal{E}$, and as long as $N(\lambda)\gg n(\lambda)$, applying $\mathcal{E}$ thus has nearly the same effect as changing each of the $n(\lambda)+1$ positions of $(x_{i},w_{i})$ independently with probability $p^{\prime}$.
@@ 5 From large-alphabet PRCs to watermarking
Acknowledgements
NG was supported by a Fannie & John Hertz Foundation Fellowship and an NSF Graduate Fellowship. AM is supported in part by a Microsoft Trustworthy AI Grant, an ONR grant and a David and Lucile Packard Fellowship. We thank Gabe Schoenbach for helpful comments on an earlier draft of this work.
@@ A.2 Additional preliminaries
for some constant $c>0$, then we could construct an adversary $\mathsf{Adv}^{\prime}$ which violates security of $F_{s}$ (namely, Definition 3.1), as follows: $\mathsf{Adv}^{\prime}$ makes $Qm$ calls to $F_{s}(\cdot)$ on random inputs, which gives it samples $(x_{j}^{(i)},F_{s}(x_{j}^{(i)})\oplus e_{j}^{(i)})$ for $j\in[m],i\in[Q]$, draws a uniformly random permutation $\pi:[N]\to[N]$ and a uniformly random string $z\in\{0,1\}^{N}$, and outputs the result of $\mathsf{Adv}$ given the $Q$ tuples $y^{(i)}:=\pi\circ(((x_{1}^{(i)},F_{s}(x_{1}^{(i)})\oplus e_{1}^{(i)}),\ldots,(x_{m}^{(i)},F_{s}(x_{m}^{(i)})\oplus e_{m}^{(i)}),r^{(i)})\oplus z)$, where $r^{(i)}$ are uniformly random strings of length $N-(n+1)m$ (for $i\in[Q]$). Note that if the $mQ$ calls made by $\mathsf{Adv}^{\prime}$ were generated by a random function, then the strings $y^{(i)}$ are uniformly random in $\{0,1\}^{N}$, conditioned on the event that $x_{j}^{(i)}$ are all distinct, for $j\in[m],i\in[Q]$. The probability that all $x_{j}^{(i)}$ are distinct is at least $1-mQ/2^{n}\geq 1-\mathsf{negl}(\lambda)$, meaning that, by Eq. 5, $\mathsf{Adv}^{\prime}$ achieves advantage $\lambda^{-c}-\mathsf{negl}(\lambda)$ on distinguishing a random element of the family $\{F_{s}\}_{s}$ from a uniformly random function, thus contradicting the security of $\mathcal{F}$. ∎
@@
Fix $\lambda\in\mathbb{N}$, and write $n=n(\lambda),\ell=\ell(\lambda),m=m(\lambda),N=N(\lambda)$. Let $\mathcal{E}$ be a $p$-weakly-substitution-bounded channel over $\{0,1\}$. Given $x_{j},e_{j}$ ($j\in[m]$) drawn as in 7 of Algorithm 1 and a uniformly random string $r\sim\operatorname{Unif}(\{0,1\}^{N-(n+1)m})$, consider the string $a:=((x_{1},F_{s}(x_{1})\oplus e_{1}),\ldots,(x_{m},F_{s}(x_{m})\oplus e_{m}),r)$, and note that the output of $\mathsf{Encode}(1^{\lambda},\mathsf{sk},\emptyset)$ given that $x_{1},\ldots,x_{m},e_{1},\ldots,e_{m},r$ are sampled is $\bar{y}:=\pi\circ(a\oplus z)$. Given $v\in\{0,1\}^{n}$, we let, with slight abuse of notation, $\mathcal{E}(v)\in\{0,1\}^{n}$ denote a random variable drawn from the eponymous distribution $\mathcal{E}(v)\in\Delta(\{0,1\}^{n})$. We have